There are two bags A and B. Bag A contains 5 red balls and 3 blue balls. Bag B contains



3 red and 5 blue balls. A fair die is rolled, if 3 or 5 turns up a ball is drawn from the bag



A else the ball is drawn from the bag B. Find the probability that a blue ball is drawn.

1
Expert's answer
2022-01-07T07:04:17-0500

Solution:

P(Blue ball) = P(Blue ball from first bag) + P(Blue ball from second bag)

"=\dfrac26\times \dfrac38+\dfrac46\times \dfrac58
\\=\dfrac18+\dfrac5{12}
\\=\dfrac{13}{24}"